40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4041     Entry time: Fri Dec 10 11:04:03 2010
Author: Osamu 
Type: Update 
Category: SUS 
Subject: ETM oplev mufunctioning 

20101208_ETMX_oplev_sum.png

 

This plot shows ETM oplev and OSEM trend for 10 hours on day before yesterday as almost the same as plot shown this entry. I reported the 10-30minites fluctuations were seen, but I noticed it comes from not suspension but from oplev power fluctuation.

After Kiwamu fixed the ETM OSEM touch yesterday afternoon, still the same trend was seen, so we had thought what we fixed was not enough. This morning I looked at the yesterday's and day before yesterday's trend and noticed the simila trend both the pit and yaw in ETM oplev but not on the OSEM trend. Kiwamu suggested me to put the oplev sum on the same plot. It was!

So, ETMX is not bad, but in fact, still alignment fluctuation exist on the cavity. ITM?

 

ELOG V3.1.3-